OpenStudy (anonymous):

n a large bag of Skittles®, each of the 5 colors (red, orange, yellow, blue, green) occurs with the same probability. You reach in and select 1 candy. Find P(yellow or orange).

OpenStudy (lgbasallote):

well it depends on the total number of skittles there are and how many individual skittle colors there are...but from this type of info given i would say P(yellow) + P(orange) would you agree @waterineyes ?

OpenStudy (ganpat):

what is probability of each of them ?? can u calculate ??

OpenStudy (ganpat):

say there are n balls for each color.. so the probability for each color = n/5n = 1/5.. so probability for yellow or orange = (1/5 + 1/5) = 2/5...